OpenStudy (anonymous):

hello i need to show one side can equal the other. work is appreciated. (1-tan Θ)/(1+tan Θ)=(cot Θ-1)/(cot Θ+1)

OpenStudy (kc_kennylau):

Convert tan to sin/cos and cot to cos/sin

OpenStudy (anonymous):

i can only convert one side not both so its alot longer than that im working on it at the same time

OpenStudy (anonymous):

i can give an example problem and solution if you need it. its all trig identity's

OpenStudy (kc_kennylau):

\[\Large\frac{1-\tan\theta}{1+\tan\theta}=\frac{1-\frac{\sin\theta}{\cos\theta}}{1+\frac{\sin\theta}{\cos\theta}}=\frac{\cos\theta-\sin\theta}{\cos\theta+\sin\theta}\]

OpenStudy (kc_kennylau):

Then divide the numerator and the denominator by \(\sin\theta\)

OpenStudy (kc_kennylau):

\[\Large=\frac{\frac{\cos\theta}{\sin\theta}-1}{\frac{\cos\theta}{\sin\theta}+1}=\frac{\cot\theta-1}{\cot\theta+1}\]

OpenStudy (kc_kennylau):

I'm using LaTeX, but you can just draw it out

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

oh wait i misunderstood what you were saying back there i think you answered it thank you

OpenStudy (anonymous):

how does (1-(sina/cosa))/(1+(sina/cosa))=(cosa-sina)/cosa+sina) btw for this i am replacing theta for a